OpenStudy (anonymous):

Integral(csc^6(x)cot(x))dx

OpenStudy (anonymous):

isnt the derivative of cosecant something like "cosecant cotangent"? or maybe its negative ?

OpenStudy (anonymous):

yeah, its negative so make \(u=\csc(x), du =-\csc(x)\cot(x)\) and get \[-\int u^5du\]

OpenStudy (anonymous):

I would start by simplifying this as much as possible. csc^6 is equal to 1/sin^6. Cotangent is equal to cos/sin. \[\int\limits \frac{ 1 }{ \sin^6 x } \times \frac{ \cos x }{ \sin x } = \int\limits \frac{ \cos x }{ \sin^7 x }\] This may or may not help you, I'm just throwing this out there.

OpenStudy (anonymous):

Satellite. This looks like where I was headed...

OpenStudy (anonymous):

And look, it's a simple u sub. This is what I get for not memorizing those weird trig derivatives

OpenStudy (anonymous):

same idea as @vinnv226

OpenStudy (anonymous):

actually @vinnv226 there is not that much to memorize if you know that the derivative of secant is secant tangent, then to take the derivative of the "co" function, replace all by "co" and stick a minus sign in front of it

OpenStudy (anonymous):

for example the derivative of tangent is secant squared, so the derivative of cotangent is minus cosecant squared

OpenStudy (anonymous):

Actually you could probably do my way and make it work. What if we make this \[\int\limits \cos x \sin^{-7} x\] Let u= sin x du=cos x dx \[\int\limits u^{-7} du\]

OpenStudy (anonymous):

yeah your way will work too for sure

OpenStudy (anonymous):

So: \[(-1/6)u^{-6} = -\sin^{-6} / 6\]

OpenStudy (anonymous):

then when you get two different answers, you can figure out why they are actually the same

OpenStudy (anonymous):

Trig integrals are always interesting

OpenStudy (anonymous):

personally, i loath integration

OpenStudy (anonymous):

And actually, thats exactly the same thing you get by doing @satellite73 s method above.

OpenStudy (anonymous):

Thanks a bunch guys! Really helped